Kellie's piggy bank contains nickels, dimes, and quarters. She has twice as many quarters as she has nickels. She has twenty-five more dimes than she has quarters. In total she has $10.00 in coins. How many of each coin does Kellie have?
Here is what I have come down with so far:
q=2n
d=q+25
n+c+q=10.00
Are any of these equations near correct?

Answer by ikleyn(52905)

.
Kellie's piggy bank contains nickels, dimes, and quarters. She has twice as many quarters as she has nickels.
She has twenty-five more dimes than she has quarters. In total she has $10.00 in coins. How many of each coin does Kellie have?
Here is what I have come down with so far:

q = 2n                  <<<---=== Correct
d = q + 25              <<<---=== Correct
n + c + q = 10.00       <<<---=== Incorrect.  It should be  5n + 10d + 25q = 1000  <<<---=== YOU COUNT CENTS ! ! !

Are any of these equations near correct?


Find my notices inside your text.